Please help me rationalize and simplify this: sqrt(3y/11x)
Answer by Alan3354(69443)

Please help me rationalize and simplify this: sqrt(3y/11x)
------------
sqrt(3y/11x) = sqrt(3y)/sqrt(11x)
Multiply NUM and DEN by DEN
= sqrt(3y)*sqrt(11x)/(sqrt(11x)*sqrt(11x))
= sqrt(33xy)/11x